# AnyGPT: Unified Multimodal LLM with Discrete Sequence Modeling

### v3

@@ 3 AnyGPT
Our interest lies in facilitating the generation of any modality to any modality with LLMs. To realize this, we propose a comprehensive framework that can be uniformly trained. As illustrated in Figure 1, this framework is composed of three main components: (11{1}1) multimodal tokenizers, (22{2}2) a multimodal language model serving as the backbone, and (33{3}3) multimodal de-tokenizers. The tokenizers transform continuous non-text modalities into discrete tokens, which are subsequently arranged into a multimodal interleaved sequence. Then the sequences are trained by the language model using the next token prediction training objective. During the inference process, multimodal tokens are decoded back into their original representations by the associated de-tokenizers. To enrich the quality of generation, multimodal enhancement modules can be deployed to post-process the generated results, including applications like voice cloning or image super-resolution. In the following section, we will introduce the details of each module.
@@ 5.1.1 Image
Method
CIDEr $\uparrow$
Flamingo (9B)  (Alayrac et al., 2022)
79.4
Flamingo (80B)
84.3
Emu (14B)  (Sun et al., 2023b)
112.4
DreamLLM (8B)  (Dong et al., 2023)
115.4
InstructBLIP (14B)   (Dai et al., 2023)
102.2
SEED-LLaMA (8B)  (Ge et al., 2023b)
123.6
AnyGPT (8B)
107.5
@@
[…]
@@
[…]
@@ 5.1.3 Music
[…]
@@
[…]
@@ Appendix D Prompts for Constructing Multimodal Interleaved Instruction Data
[…]
@@
Prompt: You are a creative assistant. I am now asking you to help me brainstorm some chatting scenarios where the user asks the agent for help. Note that the scenarios should be between ordinary people and a helpful chatbot, and it should not be an academic discussion! During the conversation, the speakers can use images or music to help convey information (but do not use video!). And the user should ask questions about it if he/she provides an image or a piece of music. Note that the image should not be a chart. Note that the images and music should not be the famous masterpieces that may arise copyright issues. Here are some of my ideas, and please show me more in the same format as mine. {demonstrations} Here’s the topics for you to try: {topics} Now it’s your turn. In these scenarios, {requirements}. GPT4: {synthetic scenarios of the provided topics, following requirements}
@@
Prompt: You are helping me to write conversations about a user talking to a chatbot named AnyGPT. In the conversations, both the user can provide images or music to help express her/his needs and ideas. And the chatbot AnyGPT can also respond to the user with images or music in its utterances. The images and music in the chat are in the form of image descriptions and music descriptions like [image: description] and [music: description], respectively. The user should provide images and music in this format and the chatbot will respond to the user like this as well. Note that at most one music appears in one conversation and the description of music should be straightforward, focusing on genres and instruments, and never mention a known music directly. Before each conversation, I will first show you a scenario and you can start writing about the chat. Here is an example: {demonstrations} Now it’s your turn for the next conversation. You only need to answer following the format in which the user and AnyGPT take turns. The conversation should be consistent with the introduction to the scenario. Remember that the utterances should be concise, try to use 5-15 words per utterance. Note that: the user utterance should always be a question or instruction. In some turns, the user provides an image or a piece of music and asks a question or makes an instruction to AnyGPT relating to the provided image or music. In other turns, the user requests AnyGPT to generate the required images or music. Note that: the description of music should focus on genres, style, and instruments. And make the description of images and music within [image: ] or [music: ] more detailed. Note that: never directly include a famous person’s name in the image descriptions or mention a piece of known music in the music description. Tips: when the user asks to convert between images and music, AnyGPT should first utter his understanding of the input image or music before generating the requested result. Keep the dialog in 2 or 3 rounds. Each dialog should include one music and at most 2 images. In this conversation, {new_scenario_description} GPT4: {A synthetic chat according to the scenario description.}

### v5

@@ 3 AnyGPT
Our interest lies in facilitating the generation of any modality to any modality with LLMs. To realize this, we propose a comprehensive framework that can be uniformly trained. As illustrated in Figure 1, this framework is composed of three main components: (1{1}) multimodal tokenizers, (2{2}) a multimodal language model serving as the backbone, and (3{3}) multimodal de-tokenizers. The tokenizers transform continuous non-text modalities into discrete tokens, which are subsequently arranged into a multimodal interleaved sequence. Then the sequences are trained by the language model using the next token prediction training objective. During the inference process, multimodal tokens are decoded back into their original representations by the associated de-tokenizers. To enrich the quality of generation, multimodal enhancement modules can be deployed to post-process the generated results, including applications like voice cloning or image super-resolution. In the following section, we will introduce the details of each module.
@@ 5.1.1 Image
Method CIDEr $\uparrow$
Flamingo (9B)  (Alayrac et al., 2022) 79.4
Flamingo (80B) 84.3
Emu (14B)  (Sun et al., 2023b) 112.4
DreamLLM (8B)  (Dong et al., 2023) 115.4
InstructBLIP (14B)   (Dai et al., 2023) 102.2
SEED-LLaMA (8B)  (Ge et al., 2023b) 123.6
AnyGPT (8B) 107.5
@@
The results of the text-to-image generation task are presented in Table 3. To ensure consistency with previous research  (Koh et al., 2023; Ge et al., 2023b; Sun et al., 2023a), we randomly select 30k images from the MS-COCO validation set and use CLIPscore as the evaluation criterion. This metric computes a similarity score between the generated image and its corresponding caption from a real image, based on CLIP-ViT-L  (Radford et al., 2021).
Method CLIPscore $\uparrow$
@@
Table 3: Comparison results on text-to-image generation task. We adopt MS-COCO captions to generate images and calculate the CLIP similarity score (CLIPscore) for evaluation.
@@ 5.1.3 Music
we evaluate AnyGPT’s performance on the MusicCaps benchmark  (Agostinelli et al., 2023) for both music understanding and generation tasks. We utilize the CLAPscore  (Wu et al., 2022; Huang et al., 2023) score as the objective metric, which measures the similarity between the generated music and a textual description.
Method CLAPscore $\uparrow$
@@
Table 6: Comparison results for music understanding and generation tasks. A metric scoring the alignment between music and textual captions is reported (CLAPscore). For music captioning, the CLAPscore of <music, real caption> pair and <music, generated caption> pair are computed for comparison.
For the evaluation of music captioning, we found that existing objective metrics may be limited in expressing the performance in the music captioning task. The diversity and subjectivity of music lead to varying opinions from individuals. Only specific music genres and instruments possess distinctive characteristics that can be easily recognized. While recent studies  (Gardner et al., 2023) have explored this issue, it remains a challenging problem to address. To ensure an objective evaluation, we compute CLAPscore of <music, real caption> pairs and <music, generated caption> pairs for comparison. These scores are averaged across the entire test set.
@@ Appendix D Prompts for Constructing Multimodal Interleaved Instruction Data
Prompt: Please list me 50 **non-academic** conversation topics about {metatopic} between an ordinary person and a helpful chatbot. Each topic should be made up of 1-10 words and the conversation contain understanding and generation of images or music. GPT4: {50 sampled topics}
Prompt: continue GPT4: {50 more sampled topics}
$\cdots\cdots$
@@
Prompt:
You are a creative assistant. I am now asking you to help me brainstorm some chatting scenarios where the user asks the agent for help.
Note that the scenarios should be between ordinary people and a helpful chatbot, and it should not be an academic discussion! During the conversation, the speakers can use images or music to help convey information (but do not use video!). And the user should ask questions about it if he/she provides an image or a piece of music. Note that the image should not be a chart. Note that the images and music should not be the famous masterpieces that may arise copyright issues. Here are some of my ideas, and please show me more in the same format as mine. {demonstrations}
Here’s the topics for you to try: {topics}
Now it’s your turn. In these scenarios, {requirements}. GPT4:
{synthetic scenarios of the provided topics, following requirements}
@@
Prompt:
You are helping me to write conversations about a user talking to a chatbot named AnyGPT.
In the conversations, both the user can provide images or music to help express her/his needs and ideas. And the chatbot AnyGPT can also respond to the user with images or music in its utterances. The images and music in the chat are in the form of image descriptions and music descriptions like [image: description] and [music: description], respectively. The user should provide images and music in this format and the chatbot will respond to the user like this as well. Note that at most one music appears in one conversation and the description of music should be straightforward, focusing on genres and instruments, and never mention a known music directly. Before each conversation, I will first show you a scenario and you can start writing about the chat. Here is an example: {demonstrations}
Now it’s your turn for the next conversation. You only need to answer following the format in which the user and AnyGPT take turns. The conversation should be consistent with the introduction to the scenario. Remember that the utterances should be concise, try to use 5-15 words per utterance. Note that: the user utterance should always be a question or instruction. In some turns, the user provides an image or a piece of music and asks a question or makes an instruction to AnyGPT relating to the provided image or music. In other turns, the user requests AnyGPT to generate the required images or music. Note that: the description of music should focus on genres, style, and instruments. And make the description of images and music within [image: ] or [music: ] more detailed. Note that: never directly include a famous person’s name in the image descriptions or mention a piece of known music in the music description. Tips: when the user asks to convert between images and music, AnyGPT should first utter his understanding of the input image or music before generating the requested result. Keep the dialog in 2 or 3 rounds. Each dialog should include one music and at most 2 images. In this conversation, {new_scenario_description}
GPT4:
{A synthetic chat according to the scenario description.}